# MAViS: A Multi-Agent Framework for Long-Sequence Video Storytelling

### v3

@@ 3.5. Character Modeling Stage
(14) $\mathcal{V}_{i}=\{G_{V}^{k}(I,p_{V}^{i}),\ G_{V}^{k}\in\mathcal{M}_{I},\ k=1,2,\dots,|\mathcal{M}_{V}|\}.$
The I2V Evaluator $E_{V}$ evaluates these candidates according to the video evaluation guide $g_{I}$, selects the best one $V_{i}$, and provides evaluation result $f^{i}_{V}$. The I2V Refiner $R_{V}$ updates the I2V prompt and produce the final video $V_{n}$ after $n$ iterations:
(15) $V_{i},f^{i}_{V}]=E_{V}(\mathcal{V}_{i},g_{V}),\quad p_{V}^{i+1}=R_{V}(f^{i}_{V},p_{I}^{i}),\quad V=V_{n}.$
Table 2. Evaluation of Automatic Metrics for Keyframe Generation and Video Animation. ”T. Flick.”, ”M. Smooth.”, ”Sub. Cons.”, ”Bg. Cons.”, ”Aesthetic”, and ”I. Quality” refer to ”Temporal Flickering”, ”Motion Smoothness”, ”Subject Consistency”, ”Background Consistency”, ‘Aesthetic Quality’, and ”Imaging Quality” from metrics in VBench (Huang et al., 2024), respectively. $\uparrow$ indicates a higher value is more desired. Bold indicates the best results.
@@
VGoT (Zheng et al., 2025) 22.37 8.53 99.00 99.31 98.94 98.74 80.11 64.59
Mora (Yuan et al., 2024) 33.98 12.68 97.32 99.23 95.23 94.88 64.36 71.48
MovieAgent (Wu et al., 2025) 31.71 10.72 97.07 99.20 93.29 94.51 63.39 71.26
@@
Table 3. User Study on the Performance of Long-Sequence Video Storytelling (Voting Results). ”Narrative”, ”Visual”, ”User Align.”, ”Sub. Cons.”, ”Sub. Natural.”, ”Bg. Cons.”, and ”Bg. Real.” refer to ”Narrative Expressiveness”, ”Visual Quality”, ”User Prompt Alignment”, ”Subject Consistency”, ”Character Naturalness”, ”Background Consistency”, and ”Background Realism”, respectively. $\uparrow$ indicates a higher value is more desired. Bold indicates the best results.
@@
VGoT (Zheng et al., 2025) 3.39 3.39 6.79 5.56 3.07 4.14 3.24
Mora (Yuan et al., 2024) 15.18 16.96 13.57 15.77 13.00 13.69 15.29
MovieAgent (Wu et al., 2025) 9.46 11.79 11.96 12.37 12.27 12.07 10.97
@@
The difference is that the I2V process incorporates additional evaluation complexity. The image evaluation guide assesses candidates based on three primary axes: 1) Visual Quality: clarity, detail resolution, and aesthetic harmony; 2) Naturalness: identifying distortions, anatomical inaccuracies, or proportion inconsistencies; and 3) Prompt Consistency: if the image content aligns with the input prompt. In contrast, the video evaluation guide extends these criteria by introducing two additional dimensions: 1) Subject Consistency: temporal stability in facial features, attire, hairstyle, and skin tone across frames; and 2) Dynamics: the richness, fluidity, and pacing of character movement throughout the clip. Together, the image and video evaluation guides work in concert to enforce the completeness of the final generation.
3.6. Generation Stages
The generation stages include: 1) the keyframe generation stage, which comprises the T2I Prompt Generator, T2I Evaluator, and T2I Refiner, responsible for generating the initial reference image for each shot (i.e., the keyframes); 2) the video animation stage, which includes the I2V Prompt Generator, I2V Evaluator, and I2V Refiner, and completes the video shot based on the keyframes; and 3) the audio generation stage, which consists of the Voice Designer, Voice Reviewer, and Subtitle Refiner, synthesizes voice-overs and background music aligned with the emotional rhythm of each shot.
@@ 3.6. Generation Stages
(16) $S_{A}^{1}=G_{A}(S),\quad f_{A}^{i}=E_{A}(S_{A}^{i}),\quad S_{A}^{i+1}=G_{A}(S,f_{A}^{i}),\quad S_{A}=S_{A}^{n}.$
@@
(17) $A^{i}_{j}=G_{A}(p_{A,j}^{i}),\quad p_{A,j}^{i+1}=R_{A}(p_{A,j}^{i};A^{i}_{j},V_{j}),\quad A_{j}=A^{n}_{j}.$
Finally, audio tracks are synchronized with video clips to produce the final long-sequence video telling with synchronized visuals and sound: $\hat{\mathcal{V}}=\{\{V_{j},\ A_{j}\}\mid j=1,2,\dots,N\}$.
4. Experiment
To evaluate the performance of MAViS, we compared it with three open-sourced long video generation frameworks and conducted fine-grained ablation studies. Further experimental details and results are provided in the supplementary material.
Implementation Details
All agents in our framework are implemented using AutoGen (Wu et al., 2024) and powered by GPT-4o (Achiam et al., 2023). The T2I model pool comprises FLUX.1 (BlackForestLabs, 2023), Stable Diffusion 3.5 (Rombach et al., 2022), and GPT-Image (OpenAI, 2025), while the I2V model pool includes Veo2 (Google, 2024), Gen-3 (Runway, 2024a), Gen-4 (Runway, 2024b), LUMA (Luma, 2024), HunyuanVideo-I2V (Kong et al., 2024), and Wan2.1 (Ang and Baole, 2025). We use MUDI (Jang et al., 2024) for LoRA training and Hailuo (MiniMax, 2024) for the T2A generation. The maximum number of 3E iterations is set to 4 for script writing, shot designing, and voice designing; 2 for the T2I process; 1 for the I2V process; and 5 for subtitle refining. As the first work targeting end-to-end long-sequence video storytelling, no suitable benchmark dataset exists. Therefore, we constructed a test set of 20 user prompts, each targeting a one-minute video for evaluation. All experiments were conducted on one NVIDIA V100 GPUs.
Metrics
We employ CLIP and Inception Score to evaluate prompt consistency and generation quality in the keyframe generation stage, and use VBench (Huang et al., 2024) to assess video animation performance. Additionally, a user study is conducted to assess real user preferences and validate the practical effectiveness of frameworks. For all evaluation metrics, higher values indicate better performance.
Baselines
We compare MAViS with three long video generation baselines: VGoT, Mora, and MovieAgent. Among them, Mora and MovieAgent adopt a modular design and are equipped with the same image and video generation models as MAViS, and are also powered by GPT-4o. Since Mora and MovieAgent do not support script generation, we provide them with scripts produced by our script writing stge. For MovieAgent, which requires manual training of LoRA models, we generate character images based on the provided scripts and train LoRA models for each script.
Table 4. Ablation study on the Structure Reviewer $E_{\mathrm{str}}$, the Content Reviewer $E_{\mathrm{con}}$, and the Style Reviewer $E_{\mathrm{sty}}$. $\uparrow$ indicates a higher value is more desired. Gray-shaded cells indicate values that are significantly lower than the others.
@@ 4. Experiment
4.1. Performance Comparisons
We evaluate each method via automatic metrics and user study. It’s worth noting that MAViS is the only framework that provides multimodal output – videos with narratives and background music.
4.1.1. Automatic Metrics
Table 2 presents the shot-level evaluation results for both keyframe generation and video animation. MAViS achieves the highest scores in CLIP (4.22) and Inception (12.81) scores, attributed to the enhanced T2I quality enabled by the 3E Principle in the T2I process. In video animation, MAViS slightly lags behind VGoT in subject consistency and background consistency but outperforms all baselines in temporal flickering, motion smoothness, and image quality. This is partly because VGoT produces videos with limited dynamics and highly saturated visuals, which improves intra-shot consistency and aesthetic quality but at the cost of realism. Overall, MAViS achieves the best comprehensive performance.
4.1.2. User Study
Table 3 summarizes the voting results from 60 evaluators across multiple aspects of comparison among the baselines. MAViS consistently outperforms all other methods in every evaluation dimension, with an average voting share of 69.44%. This highlights both the effectiveness of our architecture and the critical role of the 3E Principle. Detailed settings of the user study are provided in the supplementary material.
4.2. Ablation Study
@@
4.2.1. Script Writing Guidelines
@@
Table 5. Ablation study of the 3E Principle in the T2I and I2V Processes. ”Natural.” refers to the Naturalness score, $\uparrow$ indicates a higher value is more desired. Gray-shaded cells indicate values that are significantly lower.
@@
Table 6. Ablation study on the Shot Reviewer $E_{\textrm{shot}}$, the Voice Reviewer $E_{A}$, and the Subtitle Refiner $R_{A}$. ”C. Rate” refers to ”Compliance Rate”. $\uparrow$ indicates a higher value is more desired. Gray-shaded cells indicate values that are significantly lower.
@@
4.2.2. 3E Principle in T2I and I2V Processes
Since the evaluators and refiners in both the T2I and I2V processes cannot operate independently, we conduct an ablation study to evaluate the overall impact of the 3E Principle on these modules. The CLIP score, Inception score, Naturalness score (defined as the proportion of generated images or videos that appear physically plausible and naturally rendered), along with Subject Consistency and Dynamic Degree from the VBench metrics, are used to assess key aspects targeted by the T2I and I2V Evaluators—specifically, prompt consistency, visual quality, naturalness, subject consistency, and dynamics. As shown in Table 5, removing the 3E Principle leads to a noticeable drop in performance across both T2I and I2V outputs, particularly in terms of naturalness ($72.0\rightarrow 98.0$ in T2I and $60.0\rightarrow 92.5$ in I2V). This highlights the critical role of the 3E Principle in improving generation quality. An exception is observed in the I2V process: the Dynamic Degree is higher without the 3E Principle ($44.58>35.63$). This may be because high-dynamic videos are more prone to physical inconsistencies and instability, leading the agents to favor lower-dynamic, more stable results when guided by the 3E Principle.
4.2.3. Shot Reviewer, Voice Reviewer, and Subtitle Refiner
@@ 5. Conclution
In this work, we presented MAViS, a multi-agent end-to-end framework for long-sequence video storytelling. By orchestrating multiple stages within a unified pipeline, MAViS effectively addresses the key limitations of prior approaches — poor assistive capability, limited visual expressiveness, and incomplete outputs. Central to our framework is the 3E Principle (Explore, Examine, Enhance), which enables iterative refinement across all stages, maximizing output quality. Additionally, our proposed Script Writing Guidelines improve compatibility between narrative structure and generative tools. Experimental results validate that MAViS significantly advances the assistive capability, visual quality, and expressiveness of long-sequence video storytelling.

### v4

@@ 3.5 Character Modeling Stage
$\mathcal{V}_{i}=\{G_{V}^{k}(I,p_{V}^{i}),\ G_{V}^{k}\in\mathcal{M}_{I},\ k=1,2,\dots,|\mathcal{M}_{V}|\}.$ (14)
The I2V Evaluator $E_{V}$ evaluates these candidates according to the video evaluation guide $g_{I}$, selects the best one $V_{i}$, and provides evaluation result $f^{i}_{V}$. The I2V Refiner $R_{V}$ updates the I2V prompt and generates the final video $V_{n}$ after $n$ iterations:
$V_{i},f^{i}_{V}]=E_{V}(\mathcal{V}_{i},g_{V}),\quad p_{V}^{i+1}=R_{V}(f^{i}_{V},p_{I}^{i}),\quad V=V_{n}.$ (15)
@@
VGoT Zheng et al. (2025) 22.37 8.53 99.00 99.31 98.94 98.74 80.11 64.59
Mora Yuan et al. (2024) 33.98 12.68 97.32 99.23 95.23 94.88 64.36 71.48
MovieAgent Wu et al. (2025b) 31.71 10.72 97.07 99.20 93.29 94.51 63.39 71.26
@@
Table 2: Evaluation of Automatic Metrics for Keyframe Generation and Video Animation. "T. Flick.", "M. Smooth.", "Sub. Cons.", "Bg. Cons.", "Aesthetic", and "I. Quality" refer to "Temporal Flickering", "Motion Smoothness", "Subject Consistency", "Background Consistency", ‘Aesthetic Quality’, and "Imaging Quality" from metrics in VBench Huang et al. (2024), respectively. $\uparrow$ indicates a higher value is more desired. Bold indicates the best results.
@@
VGoT Zheng et al. (2025) 3.39 3.39 6.79 5.56 3.07 4.14 3.24
Mora Yuan et al. (2024) 15.18 16.96 13.57 15.77 13.00 13.69 15.29
MovieAgent Wu et al. (2025b) 9.46 11.79 11.96 12.37 12.27 12.07 10.97
@@
Table 3: User Study on the Performance of Long-Sequence Video Storytelling (Voting Results). "Narrative", "Visual", "User Align.", "Sub. Cons.", "Sub. Natural.", "Bg. Cons.", and "Bg. Real." refer to "Narrative Expressiveness", "Visual Quality", "User Prompt Alignment", "Subject Consistency", "Character Naturalness", "Background Consistency", and "Background Realism", respectively. $\uparrow$ indicates a higher value is more desired. Bold indicates the best results.
The difference is that the I2V process incorporates additional evaluation complexity. The image evaluation guide assesses candidates based on three primary axes: 1) Visual Quality: clarity, detail resolution, and aesthetic harmony; 2) Naturalness: identifying distortions, anatomical inaccuracies, or proportion inconsistencies; and 3) Prompt Consistency: if the image content aligns with the input prompt. In contrast, the video evaluation guide extends these criteria by introducing: 1) Subject Consistency: temporal stability in facial features, attire, hairstyle, and skin tone across frames; and 2) Dynamics: the richness, fluidity, and pacing of character movement throughout the clip. Together, the image and video evaluation guides work in concert to enforce the completeness of the final generation.
3.6 Generation Stages
The generation stages include: 1) the keyframe generation stage, which comprises the T2I Prompt Generator, T2I Evaluator, and T2I Refiner, responsible for generating the initial reference image for each shot; 2) the video animation stage, which includes the I2V Prompt Generator, I2V Evaluator, and I2V Refiner, and completes the video shot based on the keyframes; and 3) the audio generation stage, which consists of the Voice Designer, Voice Reviewer, and Subtitle Refiner, synthesizes voice-overs and background music aligned with the emotional rhythm of each shot.
@@ 3.6 Generation Stages
$S_{A}^{1}=G_{A}(S),\quad$ $f_{A}^{i}=E_{A}(S_{A}^{i}),\quad S_{A}^{i+1}=G_{A}(S,f_{A}^{i}),$
$\quad S_{A}=S_{A}^{n}.$ (16)
@@
$A^{i}_{j}=G_{A}(p_{A,j}^{i}),\ p_{A,j}^{i+1}=R_{A}(p_{A,j}^{i};A^{i}_{j},V_{j}),\ A_{j}=A^{n}_{j}.$ (17)
Finally, audio tracks are synchronized with video clips to form the final long-sequence video telling with synchronized visuals and sound: $\hat{\mathcal{V}}=\{\{V_{j},\ A_{j}\}\mid j=1,2,\dots,N\}$.
4 Experiment
To evaluate the effectiveness of MAViS, we benchmark it against three three long video generation baselines: VGoT, Mora, and MovieAgent, and conduct detailed ablation studies. We constructed a test set of 20 user prompts, each targeting a one-minute video for evaluation. We employ CLIP, Inception Score, and VBench Huang et al. (2024) for quantitative evaluation, and conduct a user study for subjective assessment. Further experimental settings, metrics, and baseline details are provided in the Appendix.
4.1 Performance Comparisons
We evaluate each method via automatic metrics and user study. It’s worth noting that MAViS is the only framework that provides multimodal output – videos with narratives and BGM.
4.1.1 Automatic Metrics
Table 2 presents the shot-level evaluation results for both keyframe generation and video animation. MAViS achieves the highest scores in CLIP (4.22) and Inception (12.81) scores, attributed to the enhanced T2I quality enabled by the 3E Principle in the T2I process. In video animation, MAViS slightly lags behind VGoT in subject consistency and background consistency but outperforms all baselines in temporal flickering, motion smoothness, and image quality. This is partly because VGoT crafts videos with limited dynamics and highly saturated visuals, which improves intra-shot consistency and aesthetic quality but at the cost of realism. Overall, MAViS achieves the best comprehensive performance.
@@ 4.1.1 Automatic Metrics
Table 4: Ablation study on the Structure Reviewer $E_{\mathrm{str}}$, the Content Reviewer $E_{\mathrm{con}}$, and the Style Reviewer $E_{\mathrm{sty}}$. $\uparrow$ indicates a higher value is more desired. Gray-shaded cells indicate values that are significantly lower than the others.
4.1.2 User Study
Table 3 summarizes the voting results from 60 evaluators across multiple aspects of comparison among the baselines. MAViS consistently outperforms all other methods in every evaluation dimension, with an average voting share of 69.44%. This highlights both the effectiveness of our architecture and the critical role of the 3E Principle.
4.2 Ablation Study
@@
4.2.1 Script Writing Guidelines
@@
Table 5: Ablation study of the 3E Principle in the T2I and I2V Processes. "Natural." refers to the Naturalness score, $\uparrow$ indicates a higher value is more desired. Gray-shaded cells indicate values that are significantly lower.
@@
Table 6: Ablation study on the Shot Reviewer $E_{\textrm{shot}}$, the Voice Reviewer $E_{A}$, and the Subtitle Refiner $R_{A}$. "C. Rate" refers to "Compliance Rate". $\uparrow$ indicates a higher value is more desired. Gray-shaded cells indicate values that are significantly lower.
4.2.2 3E Principle in T2I and I2V Processes
Since the evaluators and refiners in both the T2I and I2V processes cannot operate independently, we conduct an ablation study to evaluate the overall impact of the 3E Principle on these modules. The CLIP score, Inception score, Naturalness score (defined as the proportion of generated images or videos that appear physically plausible and naturally rendered), along with Subject Consistency and Dynamic Degree from the VBench metrics, are used to assess key aspects targeted by the T2I and I2V Evaluators—specifically, prompt consistency, visual quality, naturalness, subject consistency, and dynamics. As shown in Table 5, removing the 3E Principle leads to a noticeable drop in performance across both T2I and I2V outputs, particularly in terms of naturalness ($72.0\rightarrow 98.0$ in T2I and $60.0\rightarrow 92.5$ in I2V). This highlights the critical role of the 3E Principle in improving generation quality. An exception is observed in the I2V process: the Dynamic Degree is higher without the 3E Principle ($44.58>35.63$). This may be because high-dynamic videos are more prone to physical inconsistencies and instability, leading the agents to favor lower-dynamic, more stable results.
4.2.3 Shot Reviewer, Voice Reviewer, and Subtitle Refiner
@@ 5 Conclution
In this work, we presented MAViS, a multi-agent collaborative framework designed to assist in long-sequence video storytelling. By orchestrating multiple stages within a unified pipeline, MAViS effectively addresses the key limitations of prior approaches — poor assistive capability, limited visual expressiveness, and incomplete outputs. Central to our framework is the 3E Principle (Explore, Examine, Enhance), which enables iterative refinement across all stages, maximizing output quality. Additionally, our proposed Script Writing Guidelines improve compatibility between narrative structure and generative tools. Experimental results validate that MAViS significantly advances the assistive capability, visual quality, and expressiveness of long-sequence video storytelling.
Appendix A More Related Works
A.1 Video Generation
Long video generation He et al. (2022); Ge et al. (2022); Wang et al. (2024c) aims to produce longer videos than previous text-to-video (T2V) ModelScopeTeam (2023); Wang et al. (2024b) and image-to-video (I2V) Girdhar et al. (2023); Xing et al. (2025); Zhang et al. (2023) models. Auto-regressive approaches Weng et al. (2024); Henschel et al. (2024) extend videos by conditioning on earlier frames, yet often result in repetitive motions with limited narrative development. Attempts to directly integrate storytelling and video generation Yin et al. (2023b); Zheng et al. (2025) suffer from poor visual quality and unstructured video composition, leading to subpar viewer experiences. Meanwhile, commercial models have achieved notable success in both T2V GenmoTeam (2024); Kong et al. (2024); OpenAI (2024) and I2V generation Google (2024); Luma (2024); MiniMax (2024); Ang and Baole (2025), and some  KlingAI (2024); Runway (2024a, b) support temporal extension.
A.2 Image Generation
Recent advances in image generation Li et al. (2019); Xu et al. (2018); Zhang et al. (2017); Ramesh et al. (2021) have been propelled by deep learning techniques.To fulfill the task of visual storytelling, some works Liu et al. (2024); Yang et al. (2024b); Rahman et al. (2023) generate image sequences from text and image-caption pairs. To ensure subject consistency Gal et al. (2022), DreamBooth Ruiz et al. (2023) adopts subject-specific fine-tuning, while GPT-Image-1 OpenAI (2025) introduces context-aware generation for identity preservation. Recent models such as NanoBanana Google (2025) and HunyuanImage 3.0 Cao et al. (2025) further advance high-fidelity and instruction-following generation, offering stronger semantic control and compositional reasoning capabilities. In this work, MAViS leverages off-the-shelf models to generate keyframes for each shot, which serve as the basis for subsequent video animation.
A.3 AI Agents
AI agents, empowered by large language models Brown et al. (2020); Touvron et al. (2023a, b), have shown broad applicability in reasoning, planning, web interaction, and visual generation. General-purpose LLMs underpin agent frameworks Yang et al. (2023a); Yin et al. (2023a); Li et al. (2023); Hong et al. (2023); Park et al. (2023), while multimodal models Achiam et al. (2023); Anil et al. (2023) extend agent capabilities across modalities Yang et al. (2023b); Zhang and Zhang (2024); Zheng et al. (2024); Wang et al. (2024e); Liu et al. (2025); Wang et al. (2024d); Hang et al. (2024); Liao et al. (2025). In video generation, systems Yang et al. (2024a); Yuan et al. (2024); Wang et al. (2024a) explore world modeling, multi-agent coordination, and retrieval-augmented synthesis, yet often yield static content with limited transitions. Hierarchical pipelines Xie et al. (2024); Wu et al. (2025b) introduce script-to-shot workflows but still rely on laborious scripting and LoRA tuning.
Appendix B Explanation and Discuss about the Script Writing Guideline
Due to space limitations in the main text, we provided only a brief overview of the Script Writing Guideline. In this section, we offer a more detailed explanation of the guideline and elaborate on the rationale behind each design constraint.
In the Script Writing Guideline, "Scriptwriters should minimize transitions between tightly connected spatial locations" refers to avoiding consecutive shots set in spatially adjacent backgrounds—for example, the first shot showing the character entering a room by going through a door, and the next shot showing the character from the front inside the room. Such transitions can break immersion due to background inconsistencies, such as mismatched door appearances between shots.
The script writing guidelines are based on practical experience to improve compatibility with generation tools. E.g., "not to place the same character in consecutive shots" stems from the fact that such sequences often involve multi-angle cuts in the same setting. Given current model limitations in maintaining background consistency, this approach is not yet reliably achievable.
The constraints in the script writing guideline are based on the following empirical observations:
$\bullet$ Avoid placing successive shots in the same background: We measured VBench background consistency for adjacent shots sharing the same background and found it to be only 78.42, compared to 95.88 of single-shot. This justifies avoiding such situations.
$\bullet$ Minimize transitions between tightly connected spatial locations: For pairs of shots with strong spatial connections, we used GPT to assess whether their backgrounds were spatially related as described in the script. Only 26 out of 100 pairs were judged to be consistent, indicating such transitions should be avoided.
$\bullet$ Avoid complex actions within a single shot: Shots with complex actions achieved a VBench overall consistency of 23.25, compared to 27.76 for simple actions, suggesting simpler actions improve generation reliability.
$\bullet$ Avoid fine-grained visual elements within a single shot: Shots containing fine-grained visual elements scored 22.68 in VBench overall consistency, while those without such elements scored 27.76. Thus, they should be avoided.
$\bullet$ Align style with user’s quest: Using GPT to judge alignment with user instructions, only 82 out of 100 scripts were fully aligned without constraints, compared to 100 with constraints in place.
$\bullet$ Ensure script structural completeness: A complete structure is essential for enabling subsequent stages of the generation pipeline.
For more details, refer to the Script Writing Guideline, please refer to the system messages.
To evaluate the performance of MAViS, we compared it with three open-sourced long video generation frameworks. Further experimental details and results are provided in the Appendix.
[…]
DreamFactory $7.03^{{\dagger}}$ $169.71^{{\dagger}}$ $30.92^{{\dagger}}$
MAViS - 12.81 34.22
Table 7: Compare with DreamFactory Xie et al. (2024). $\dagger$} denotes the reported score in the paper of the baseline method.
[…]
MovieDreamer $19.52^{\dagger}$ $8.64^{\dagger}$ $6.05^{\dagger}$ $0.70^{\dagger}$
MAViS 34.22 12.81 7.25 0.71
Table 8: Compare with MovieDreamer Zhao et al. (2024). $\dagger$} denotes the reported score in the paper of the baseline method. "AS" denotes Aesthetic Score used in Schuhmann et al. (2022).
[…]
LCT $60.79^{\dagger}$ $67.44^{\dagger}$ $95.65^{\dagger}$ $30.14^{\dagger}$
MAViS 63.17 72.91 95.92 34.22
Table 9: Compare with LCT Guo et al. (2025). $\dagger$} denotes the reported score in the paper of the baseline method. “Consistency (avg.)” represents the average score of subject and background consistency. "P. Consistency" denotes Prompt Consistency.
Appendix C Implementation Details
All agents in our framework are implemented using AutoGen Wu et al. (2024) and powered by GPT-4o Achiam et al. (2023). The T2I model pool comprises FLUX.1 BlackForestLabs (2023), Stable Diffusion 3.5 Rombach et al. (2022), and GPT-Image OpenAI (2025), while the I2V model pool includes Veo2 Google (2024), Gen-3 Runway (2024a), Gen-4 Runway (2024b), LUMA Luma (2024), HunyuanVideo-I2V Kong et al. (2024), and Wan2.1 Ang and Baole (2025). We use MUDI Jang et al. (2024) for LoRA training and Hailuo MiniMax (2024) for the T2A generation. The maximum number of 3E iterations is set to 4 for script writing, shot designing, and voice designing; 2 for the T2I process; 1 for the I2V process; and 5 for subtitle refining. As the first work targeting the efficient translation of single lines of ideas to long-sequence video storytelling, no suitable benchmark dataset exists. Therefore, we constructed a test set of 20 user prompts, each targeting a one-minute video for evaluation.
Metric Question
Narrative Expressiveness Which video performed best overall in terms of narrative engagement, story coherence, and viewing experience?
Visual Quality Which video was the most visually expressive in terms of emotion, atmosphere, and cinematic feel?
User Prompt Alignment Which video best aligns with the original user prompt (e.g., character setup, plot elements, or keywords)?
Character Consistency Which video had the most consistent character appearances across scenes? (clothing, hairstyle, facial features)
Character Naturalness Which video had the most natural and correct character generation? Consider anatomical realism, physical motion plausibility,
and the absence of duplicate or mistakenly generated characters.
Background Consistency Which video had the most consistent and logically coherent background style across shots?
Background Realism Which video had the most natural background elements and physically plausible movements?
Table 10: Metrics and Questions for User Study.
Method Keyframe Generation Video Animation
VGoT Zheng et al. (2025) 22.32 8.41 99.00 99.28 98.98 98.76 79.95 64.62
Mora Yuan et al. (2024) 34.02 12.70 97.25 99.21 94.58 95.01 64.41 71.39
MovieAgent Wu et al. (2025b) 31.64 10.69 97.05 99.18 93.44 94.62 63.38 91.25
MAViS (ours) 34.15 12.79 99.08 99.52 95.54 96.10 64.05 72.68
Table 11: Compare with baselines on MovieBench and ViStoryBench.
To standardize intermediate outputs for downstream processes, we introduced a Script Consolidator agent in the Script Writing stage. This agent formats the script after each 3E iteration to ensure consistency. In both the T2I and I2V processes, due to limitations on the number of images that can be processed simultaneously, the T2I Evaluator and I2V Evaluator adopt a batched evaluation strategy. Following this, T2I Selector and I2V Selector agents are employed to select the best candidates based on all evaluation results. Specifically, the I2V Evaluator samples 8 evenly sampled frames from each video candidate for evaluation.
The T2I process in the Shot Designing stage mirrors that of Character Modeling, with the difference being that the T2I model pool in the Shot Designing stage incorporates the LoRA models trained during Character Modeling. We use MUDI Jang et al. (2024) for LoRA training, which supports multi-subject generation, allowing a single LoRA to serve an entire script.
[…]
The T2I model pool includes the open-source models FLUX BlackForestLabs (2023) and Stable Diffusion 3.5 StabilityAI (2024), as well as the API-based GPT-Image OpenAI (2025). The I2V model pool comprises open-source models Hunyuan Kong et al. (2024) and Wan2.1 Ang and Baole (2025), along with API-based models Gen3 Runway (2024a), Gen4 Runway (2024b), LUMA Luma (2024), and Veo2 Google (2024). For T2A, we use the Minimax Hailuo API MiniMax (2024). All experiments were conducted on one NVIDIA V100 GPUs.
More samples generated by MAViS framework are shown in Figure. 6 and Figure. 7.
C.1 Metrics
we constructed a test set of 20 user prompts, each targeting a one-minute video for evaluation. We employ CLIP and Inception Score to evaluate prompt consistency and generation quality in the keyframe generation stage, and use VBench Huang et al. (2024) to assess video animation performance. Additionally, a user study is conducted to assess real user preferences and validate the practical effectiveness of MAViS and baselines. For all evaluation metrics, higher values indicate better performance.
C.2 Baselines
We compare MAViS with three long video generation baselines: VGoT, Mora, and MovieAgent. Among them, Mora and MovieAgent adopt a modular design and are equipped with the same image and video generation models as MAViS, and are also powered by GPT-4o. Since Mora and MovieAgent do not support script generation, we provide them with scripts formed by our script writing stge. For MovieAgent, which requires manual training of LoRA models, we generate character images based on the provided scripts and train LoRA models for each script.
Appendix D Compare with More Baselines
In our experiments, we conducted direct comparisons with three open-source baselines: VGoT Zheng et al. (2025), Mora Yuan et al. (2024), and MovieAgent Wu et al. (2025b). Subsequently, we compared our results with the reported performances of non-open-source baselines.
We first compare our method with DreamFactory Xie et al. (2024). Given the scarcity of real videos in long-sequence video storytelling with assistive generation, computing the FID metric is not feasible. As shown in Table 7, MAViS achieves higher scores across most metrics; however, there is a substantial gap in Inception Score. We adopted a widely used implementation of the Inception Score, which may differ from the version used by DreamFactory, potentially leading to differences in score range. Since the baseline is not open-sourced, we are unable to verify the exact cause of this discrepancy.
Next, we compare our method with MovieDreamer Zhao et al. (2024). As shown in Table 8, MAViS outperforms this baseline across all evaluation metrics.
Subsequently, we compare our method with LCT Guo et al. (2025), using evaluation metrics sourced entirely from VBench. As shown in Table 9, MAViS consistently outperforms this baseline across all metrics.
Appendix E User Study
In the User Study, participants were asked to vote for the best video in each group based on seven evaluation criteria. As listed in Table 10, the criteria include: Narrative Expressiveness, Visual Quality, User Prompt Alignment, Character Consistency, Character Naturalness, Background Consistency, and Background Realism. Prior to voting, detailed guidelines for each metric were provided to ensure consistency and reliability in evaluation. Some screenshots of the user study HTML interface is shown in Figures 3,  4, and 5.
Appendix F Evaluation on the Existing Dataset
We are establishing an efficient agentic system from users’ single line of ideas to final pixels plus other multimodalities. However, existing long video benchmarks, MovieBench Wu et al. (2025a) and ViStoryBench Zhuang et al. (2025), focus only on benchmarking T2I and I2V models’, not on the entire agentic system. In particular, they ignore the script generation and scripts’ compatibility with existing T2I and I2V models’ capabilities.
Nevertheless, we sampled three examples (each consisting of 10 shots) from both MovieBench and ViStoryBench. As shown in Table 11, MAViS consistently outperforms the compared methods.
Appendix G Quality-Efficiency Analysis
In the stages where the 3E principle is applied, we evaluated the impact of the maximum iteration number ($I$) on generation quality and time using the same metrics as in our paper.It’s worth noting that C. Rate refers to Compliance Rate, defined as the proportion of outputs that fully adhere to their respective guidelines. It is used to measure the success rate of each stage’s generation.
II Narrative Visual User Align. C.Rate Time(Second)
0 12.58 18.44 19.56 74.00 9.80
1 25.32 22.85 24.40 86.00 20.90
2 32.54 24.82 26.49 94.50 28.24
3 33.48 25.60 26.52 98.50 36.13
4 34.00 26.00 26.67 100.00 41.58
Table 12: Quality-Efficiency Analysis on scriptwriting.
For script writing, as shown in Tabel 12, the performance is already satisfactory at $I=2$. Increasing $I$ further can improve generation quality with minimal impact on overall efficiency.
II CLIP Inception Natural. C.Rate Time(Second)
0 31.25 10.54 72.00 68.00 175.19
1 32.25 11.30 84.00 81.00 381.32
2 34.22 12.81 98.00 92.50 557.47
3 34.26 12.78 98.50 96.00 602.06
Table 13: Quality-Efficiency Analysis on Keyframe Generation.
II CLIP Inception Natural. Sub. Cons. Dyn. Deg. C.Rate Time(Minute)
0 32.05 11.22 60.00 93.00 44.58 72.50 61.17
1 34.53 13.22 92.50 95.72 35.62 86.00 124.21
2 34.62 13.31 95.00 95.51 35.68 94.00 176.77
3 34.35 13.18 96.50 95.46 35.54 98.00 215.38
Table 14: Quality-Efficiency Analysis on Video Animation.
Shot Designing Voice Designing Subtitle Refining
II C.Rate Time(Second) C.Rate Time(Second) C.Rate Time(Second)
0 84.50 24.34 92.00 5.62 56.50 1.58
1 89.00 73.49 95.00 13.48 86.00 2.83
2 94.00 122.352 98.50 26.19 92.50 3.95
3 97.00 161.21 97.00 161.21 98.00 4.36
4 98.50 195.86 99.00 29.84 99.50 4.84
5 100.00 231.60 100.00 31.68 100.00 4.89
Table 15: Quality-Efficiency Analysis on shot designing, voice designing, and subtitle refining.
For keyframe generation, as shown in Tabel 13, the performance is already satisfactory at $I=2$. Increasing $I$ further can enhance generation quality, with only a slight impact on overall efficiency.
For video animation, as shown in Tabel 14, the performance is already satisfactory at $I=2$. While increasing $I$ can further improve quality, it has a significant impact on overall efficiency. Therefore, we set $I=1$ for video animation in our experiments.
For shot designing, voice designing, and subtitle refining, as shown in Tabel 15, increasing $I$ improves generation quality with minimal impact on overall efficiency. However, since errors in these stages can critically affect the entire generation process, we set $I=5$ for these tasks in our experiments.
Additionally, since iterations may terminate before reaching the maximum iteration number, the generation time does not scale linearly with $I$.
Figure 3: Screenshot of the user study HTML interface with instructions and generated video candidates.
Figure 4: Screenshot of the user study HTML interface with seven evaluation questions.
Figure 5: Screenshot of the user study HTML interface with generated video candidates and seven evaluation questions.
Figure 6: More samples generated by MAViS framework.
Figure 7: More samples generated by MAViS framework.